y¨ + 4y´ + 4y=(3 + x)e-2x , y(0)=2, y´(0)=5


1
Expert's answer
2021-10-22T11:06:12-0400

Let us solve the differential equation "y'' + 4y' + 4y=(3 + x)e^{-2x} ,\ y(0)=2, y'(0)=5".

The characteristic equation "k^2+4k+4=0" is equivalent to "(k+2)^2=0," and hence has the solutions "k_1=k_2=-2." Therefore, the general solution of the differential equation "y'' + 4y' + 4y=(3 + x)e^{-2x}" is "y=(C_1+C_2x)e^{-2x}+y_p," where "y_p=x^2(ax+b)e^{-2x}=(ax^3+bx^2)e^{-2x}."

It follows that

"y_p'=(3ax^2+2bx)e^{-2x}-2(ax^3+bx^2)e^{-2x}\\=(3ax^2+2bx-2ax^3-2bx^2)e^{-2x},"

"y''=(6ax+2b-6ax^2-4bx)e^{-2x}-2(3ax^2+2bx-2ax^3-2bx^2)e^{-2x}\\=
(6ax+2b-6ax^2-4bx-6ax^2-4bx+4ax^3+4bx^2)e^{-2x}."

Then we have

"(6ax+2b-6ax^2-4bx-6ax^2-4bx+4ax^3+4bx^2)e^{-2x}+4(3ax^2+2bx-2ax^3-2bx^2)e^{-2x}+4(ax^3+bx^2)e^{-2x}=\n(3 + x)e^{-2x},"

and hence "6ax+2b=3 + x." We conclude that "6a=1,\ 2b=3," and thus "a=\frac{1}6,\ b=\frac{3}2."

Consequently, the general solution is of the form

"y=(C_1+C_2x)e^{-2x}+(\frac{1}6x^3+\frac{3}2x^2)e^{-2x}=(C_1+C_2x+\frac{3}2x^2+\frac{1}6x^3)e^{-2x}."

Then "2=y(0)=C_1."

It follows that "y'=(C_2+3x+\frac{1}2x^2)e^{-2x}-2(C_1+C_2x+\frac{3}2x^2+\frac{1}6x^3)e^{-2x}," and hence "5=y'(0)=C_2-2C_1=C_2-4." Then "C_2=9."


We conclude that the solution of the differential equation "y'' + 4y' + 4y=(3 + x)e^{-2x} ,\ y(0)=2, y'(0)=5" is

"y=(2+9x+\frac{3}2x^2+\frac{1}6x^3)e^{-2x}."